OpenStudy (anonymous):

write the equation of a line perpendicular to y=5x+3 and whose y intercept is 4. show all your work HELPPPPP PLEASE !!

OpenStudy (anonymous):

the slope of the line = -1/5 y = (-1/5) x + c c = 4 so equation is y = (-1/5) x + 4